# Metric Oja depth, new statistical tool for estimating the most central objects

### v2

@@ 4.1 In-sample optimization
Throughout, we abbreviate the depth functions as follows: MOD3 = metric Oja depth, MOD2 = the depth-like counterpart of MOD3 defined in Section 3.3, MHD = metric half-space depth, MLD = metric lens depth, MSD = metric spatial depth. Depths were implemented in Rcpp [11], and are available at https://github.com/vidazamani/Depth-functions-for-Object-Data. All these functions have computational complexity $\mathcal{O}(n^{3})$ for computing the depths of the full sample (assuming [7, Algorithm 1] with the sample itself as anchors is used for MHD), except for MOD3 which is instead $\mathcal{O}(n^{4})$. Note that the complexities of the Euclidean counterparts of these, and other Euclidean depths can be found in [20, Table 1.9].
In our first experiment, we compare the depths with simulations in the context of location estimation. Given a sample $X_{1},\ldots,X_{n}\in\mathcal{X}$ from a distribution $P$, each sample depth function $D_{n}$ is used to determine an in-sample estimate of the deepest object, i.e., $\hat{\mu}:=X_{i_{0}}$ where $i_{0}=\mathrm{argmin}_{i=1,\ldots,n}D_{n}(X_{i})$. The depths are compared based on the averages, over 200 replications, of the estimation errors $d(\hat{\mu},\mu)$ where $\mu\in\mathcal{X}$ is the true location of the distribution $P$. In addition to the error, we evaluate also the computation times of the depths.
@@ 4.1.1 Correlation matrix dataset
In our first simulation, we consider the Riemannian manifold $(\mathcal{X},d)$ of $p\times p$ correlation matrices as the data space, where $d(X_{1},X_{2})=\|\mathrm{Log}(X_{1}^{-1/2}X_{2}X_{1}^{-1/2})\|_{F}$ is the affine invariant Riemannian metric, see [2]. Such data are commonly encountered, e.g., in neuroscience in the form of brain connectivity matrices [23].
@@
The parameter $\nu$ was determined as follows: each $X_{i}$ had the probability $\varepsilon>0$ of being an outlier, in which case $\nu=3$, whereas the remaining $1-\varepsilon$ proportion of the data (the bulk) used $\nu=0$. The matrices $U_{i}$ are uniformly random, see [24] for details. This choice ensures that the eigenvectors of $S_{i}$ do not favor any direction. Consequently, the distribution of $S_{i}$ is rotationally symmetric around the half-line $\{CI_{p}\mid C\geq 0\}$ and, intuitively, the most central matrix w.r.t. this distribution has to reside on this half-line. As the correlation matrices $X_{i}$ are obtained from the $S_{i}$ by scaling the diagonal elements to unity, any reasonable depth measure should estimate the central object of the distribution of $X_{i}$ to be $\mu=I_{p}$. Note that both the bulk and the outliers share the same true most central object (identity matrix) but the high variance of the outliers makes the estimation more difficult for larger $\varepsilon$.
@@
The simulation has three parameters, the contamination proportion $\varepsilon=0.05,0.10$, dimension $p=3,5,10$ and sample size $n=10,20,30,40,50,60$, leading to a total of 36 cases. Note that $n$ was deliberately chosen as low, since object data scenarios typically have much smaller sample sizes than in Euclidean data analysis. The average estimation errors are shown in Figure 1, grouped by $\varepsilon$ and $p$. In all cases, as $n$ increases, the estimation error gradually decreases. Moreover, the complexity of the data space grows as the dimension $p$ of the matrices increases, leading to a decline in accuracy. Increasing the percentage of outliers from $5\%$ to $10\%$ did not lead to significant changes in the results, showing that the performance of the estimators remains stable and indicating that they are indeed robust. In all cases, MOD3 and MOD2 demonstrated the best performance, respectively, while MHD showed the weakest performance. MSD was the third most efficient method and close to MOD3 and MOD2 in performance. From the timing results given in Figure 2, we observe that neither $p$ nor $\varepsilon$ affect the running time. This is because neither has an effect on the size of the sample distance matrix, see also the discussion in Section 6. Additionally, MOD3 and MOD2 have a relatively high time cost, see Figure 2, meaning that these depths essentially offer improved performance at the cost of speed. An alternative version of Figure 1 where the results are shown relative to MSD is presented in Appendix D.
@@ 4.1.2 Hypersphere dataset
In this experiment, we generated samples of points on a $p$-dimensional unit hypersphere $\mathcal{X}$. As a metric $d$ we used the usual arc length distance. Each point $X_{i}$ was simulated by generating the $p$-dimensional random vector $Z_{i}$ from $\mathcal{N}_{p}(\lambda 1_{p},I_{p})$, where $1_{p}$ is a vector of ones and $\lambda\neq 0$, and then taking $X_{i}:=Z_{i}/\|Z_{i}\|$. By symmetry, this strategy leads to the most central object always being equal to $\mu=\mathrm{sign}(\lambda)(1/\sqrt{p})1_{p}$ and the absolute value of the parameter $\lambda$ controls the spread of the points; the larger the value of $|\lambda|$, the more closely the points are concentrated around the location. For the outliers we used $\lambda=-1$ and for the bulk $\lambda=5$, meaning that the two distributions have their most central points on the opposite sides of the hypersphere, see Figure D.1 in Appendix D for an illustration.
@@
The simulation thus has three parameters, $\varepsilon$, $p$, $n$, and we use for them the same values as in the earlier simulation. As shown in Figure 3 and its relative version presented in Appendix D, the performance of each metric depth function on the hypersphere dataset is very similar to their performance on the correlation data. We have omitted the timing results for the sphere simulation as they were visually almost identical to Figure 2.
We also performed another simulation in the same setting with the parameters $\varepsilon=0.10$, $n=100,200$, $p=50,100,200,400,800,1600$, where the purpose is to investigate the effect of increasing dimension on the estimate. The average estimation errors over 100 replicates are plotted in Figure 4. From the results we observe that: (i) MSD, MLD, MOD2, MOD3 yield almost identical results and, as expected, improve when $n$ is increased. (ii) MHD is worse than the other four and changing $n$ does seems not affect it. The standard half-space depth is known to degenerate in higher dimensions [10] and, thus, it could be that MHD suffers from the same. A non-trivial solution could be to regularize its computation in a suitable way. (iii) For all methods, the error grows in $p$. From the viewpoint of efficiency, e.g., $n=100,p=500$ yields roughly the same accuracy as $n=200,p=650$. (iv) To obtain a baseline, we considered also an estimator which picks a random object from the sample and uses it as the estimate of the location. This leads, regardless of $p$, to estimation errors of $0.413$, $0.4078$ for $n=100,200$, respectively, showing that all estimators improve considerably upon this naive benchmark.
@@ 4.2 Out-of-sample optimization
The in-sample optimization studied in the previous section might come across as a naive approach, but in many metric spaces it is the best one can do, the lack of Euclidean structure preventing the use of standard optimization algorithms. However, despite being non-Euclidean, some metric spaces admit coordinate representations/encodings in Euclidean spaces, e.g., unit spheres (stereographic projection), the positive definite manifold (Cholesky decomposition) and the space of $L_{2}$-functions (Karhunen-Loève expansion, assuming we truncate it). In these cases, the full Euclidean non-linear optimization machinery can be brought forth to find the deepest point.
To formalize the earlier, assume that there exists a bijective map $f:\mathcal{X}\to\mathcal{S}$ where $S\subseteq\mathbb{R}^{q}$ for some $q$. Then, the problem of finding the deepest object of a sample $X_{1},\ldots,X_{n}\in\mathcal{X}$ can be formulated as
@@
where $D_{n}$ is some metric depth function w.r.t. the sample and $v$ is an Euclidean vector. After finding the optimizer $v_{0}$, it can be mapped to $\mathcal{X}$ simply as $f(v_{0})$. Since the map $v\mapsto D_{n}(f^{-1}(v))$ is highly non-linear, we propose using bounded Nelder-Mead (NMKB) and Limited-memory BFGS (L-BFGS-B) for the purpose. NMKB is a robust and derivative-free optimizer, which makes it generally well-suited for non-smooth functions like metric depths. However, it can be slow in high-dimensional problems. On the other hand, L-BFGS-B uses approximated gradients to find the optimum solutions and stores only the most recent steps during the optimization process, making it faster and more memory-efficient. See [4] for more details.
@@
In this simulation, the matrix dimension $p$ was set to 5, and the contamination proportion $\varepsilon$ was fixed to 10%. The sample size, as in previous simulations in Sections 4.1.1 and 4.1.2, was considered to be $n=10,20,30,40,50,60$. A new parameter introduced in this simulation was the number of principal components, determined by a threshold $\texttt{TSH}\in(0,1]$, representing the minimum proportion of total variance to be explained by the chosen principal components. Thus, the number of principal components is $r=\max(2,\{k:\sum_{i=1}^{k}\text{ExplainedVariance}_{i}\geq\text{TSH}\})$. In this simulation TSH was set to be $0.9$. The whole simulation was repeated 500 times using both MHD and MOD3. As further competitors we took the MHD-based in-sample and out-of-sample estimators in the R-package MHD [6], see also [7]. The latter is based on Riemannian optimization of the depth function and denoted by DLP hereafter. Finally, also the in-sample version of MOD3 was included.
Figure 5: Comparing the performance of different optimization methods on MOD3 and MHD. DLP refers to the out-of sample method Left: estimation error of 7 different cases. L-FBSG-B yielded exactly the same result on MOD3 as NMKB (light red line). Right: Estimation Time. The scale of the y-axis is logarithmic.
As observed in the results in Figure 5 (left), the use of the two introduced optimization algorithms (NMKB and L-BFGS-B) combined with dimension reduction has significantly contributed to reducing the estimation error compared to using only the in-sample estimator, which are indicated by the solid lines. What is conspicuous is that, NMKB has the best performance over the two other optimization strategies, DLP and L-BFGS-B. Moreover, MOD3 exhibits lower estimation error than MHD in both out-of-sample and in-sample cases. However, as shown in the Figure 5 (right), MOD3 involves a more time-consuming process compared to MHD. Additionally, despite the substantial computational burden imposed by using NMKB algorithm, we observe that with increasing sample size, even without using NMKB, the in-sample estimators tends to follow a similar trend.
@@ 5 Real data example
In this section, we aim to evaluate the performance of the proposed depth functions using real data. The secondary aim of this experiment is to demonstrate how the depth functions, whose analytically complex form prevents theoretical inferential results, can still be used for statistical inference via relying on computational techniques. For this purpose, the Age_pyramids_2014 data from R-package HistDAWass was chosen. This dataset contains histogram-valued data representing the age distribution for each country in the world, for both sexes combined and separately for males and females. We model the histograms as objects in the Wasserstein space $(\mathcal{X},d)$ where $d$ is the $L_{2}$-Wasserstein distance, implemented in the HistDAWass-package.
The current analysis was conducted on a subset of data containing both sexes and all African and European countries, accounting for a total of 95 countries. Based on the geographic locations of the two continents, we hypothesize that age distributions differ between them. This is visually confirmed in Figure 6, where the deepest age distributions, w.r.t. MOD3, among African countries (left) shows a different structure from the corresponding country in Europe (right). Thus, our research question is: Is there a significant difference in the age distribution between these two groups? To study this, we next conduct a permutation test for determining whether the difference in the deepest histograms in the two continents is statistically significant. Letting $f_{1},f_{2}\in\mathcal{X}$ denote the deepest in-sample histograms of the two groups, we use the $L_{2}$-Wasserstein distance between them, $t=d(f_{1},f_{2})$, as the test statistic. To simulate the distribution of $t$ under the null hypothesis (i.e., assuming no difference between the two groups), we randomly permute the group labels of the $n=95$ histograms and compute the test statistic value. This reshuffling was done a total of 500 times, yielding $t^{*}_{1},\ldots,t^{*}_{500}$, using which the $p$-value of the test is computed as $\{t\leq t^{*}_{i}\}/500$.
The resulting $p$-values for the five different depths are extremely small and are shown in Figure 7 (left). All depth functions agree that there is a difference between the groups at the significance level $0.05$.
@@
Figure 7: The $p$-value results of permutation test across depth functions. The test is at level $\alpha$ = 0.05, indicated by the horizontal blue dashed-line.
@@
We next take the previous result of no differences as a ground truth and continue the experiment by (a) contaminating the original data set, and (b) performing the same test for the contaminated data set and observing which of the depth-metric combinations still let us obtain the correct conclusion (no group difference). This experiment therefore mimics the practical scenario where the data has been partially wrongly recorded. To perform the contamination, we randomly select $k$ countries and swap their labels (a European country is labeled an African country and vice versa), using two different choices, $k=12,16$. To reduce randomness in the choice of the swap, we repeat the simulation 10 times, using 100 iterations in the test each time and the resulting average $p$-values are shown in Figure 7 (middle and right). Under medium contamination (12 out of 95 countries mislabeled), all depths continue to yield the correct decision. Whereas, under high contamination (16 out of 95 countries mislabeled) both MHD and MOD2 identify the groups as indistinct. The reason for this is likely that the algorithm for MHD is approximative [7] and MOD2 is not a “true” depth function and does not necessarily measure centrality of its input object, see the discussion after formula (3). The most consistent performance across swaps is given by MOD3 and MSD, which likely relates to the fact that they both use numerical distance information in their computation, unlike MHD and MLD which only use indicators, see Section 6 for more discussion of this. The results also align with the simulation studies in Section 4.1 where MSD was always almost as good as MOD3.

### v3

@@ 4.1 In-sample optimization
Throughout the paper, we use the following abbreviations for the depth functions: MOD3 (metric Oja depth), MOD2 (the depth-like counterpart of MOD3 defined in Section 3.3), MHD (metric half-space depth), MLD (metric lens depth), and MSD (metric spatial depth). The depths were implemented in Rcpp [12] and are available at https://github.com/vidazamani/Depth-functions-for-Object-Data.
In our first experiment, we compare the depths with simulations in the context of location estimation. Given a sample $X_{1},\ldots,X_{n}\in\mathcal{X}$ from a distribution $P$, each sample depth function $D_{n}$ is used to determine an in-sample estimate of the deepest object, i.e., $\hat{\mu}:=X_{i_{0}}$ where $i_{0}=\mathrm{argmin}_{i=1,\ldots,n}D_{n}(X_{i})$. The depths are compared based on the averages, over 200 replications, of the estimation errors $d(\hat{\mu},\mu)$ where $\mu\in\mathcal{X}$ is the true location of the distribution $P$. In addition to the error, we also evaluate the computation times of the depths. When measuring the running times, we disregard the computation of the interobject distances, as this step is common to all competing methods.
@@ 4.1.1 Correlation matrix dataset
In our first simulation, we consider the Riemannian manifold $(\mathcal{X},d)$ of $p\times p$ correlation matrices as the data space, where $d(X_{1},X_{2})=\|\mathrm{Log}(X_{1}^{-1/2}X_{2}X_{1}^{-1/2})\|_{F}$ is the affine invariant Riemannian metric; see [2]. Such data are commonly encountered, e.g., in neuroscience in the form of brain connectivity matrices [32].
@@
The parameter $\nu$ was determined as follows: each $X_{i}$ had the probability $\varepsilon>0$ of being an outlier, in which case $\nu=3$, whereas the remaining $1-\varepsilon$ proportion of the data (the bulk) used $\nu=0$. The matrices $U_{i}$ are uniformly random; see [33] for details. This choice ensures that the eigenvectors of $S_{i}$ do not favor any direction. Consequently, the distribution of $S_{i}$ is rotationally symmetric around the half-line $\{CI_{p}\mid C\geq 0\}$ and, intuitively, the most central matrix w.r.t. this distribution has to reside on this half-line. As the correlation matrices $X_{i}$ are obtained from the $S_{i}$ by scaling the diagonal elements to unity, any reasonable depth measure should estimate the central object of the distribution of $X_{i}$ to be $\mu=I_{p}$. Note that both the bulk and the outliers share the same true most central object (identity matrix) but the high variance of the outliers makes the estimation more difficult for larger $\varepsilon$.
@@
The simulation has three parameters, the contamination proportion $\varepsilon=0.05,0.10$, dimension $p=3,5,10$ and sample size $n=10,20,30,40,50,60$, leading to a total of 36 cases. Note that $n$ was deliberately chosen as low, since object data scenarios typically have much smaller sample sizes than in Euclidean data analysis. The average estimation errors are shown in Figure 1, grouped by $\varepsilon$ and $p$. In all cases, as $n$ increases, the estimation error gradually decreases. Moreover, the complexity of the data space grows as the dimension $p$ of the matrices increases, leading to a decline in accuracy. Increasing the percentage of outliers from $5\%$ to $10\%$ did not lead to significant changes in the results, indicating the stability and robustness of the estimators. In all cases, MOD3 and MOD2 demonstrated the best performance, respectively, while MHD showed the weakest performance. MSD was the third most efficient method and close to MOD3 and MOD2 in performance. From the timing results given in Figure 2, we observe that neither $p$ nor $\varepsilon$ affect the running time. This is because neither has an effect on the size of the distance matrix; see the discussion in Section 3.2. Additionally, MOD3 and MOD2 have a relatively high time cost; see Figure 2, meaning that these depths essentially offer improved performance at the cost of speed. An alternative version of Figure 1 where the results are shown relative to MSD is presented in Appendix C.
@@ 4.1.2 Hypersphere dataset
In this experiment, we generated samples of points on a $p$-dimensional unit hypersphere $\mathcal{X}$. As a metric $d$, we used the usual arc length distance. Each point $X_{i}$ was simulated by first generating the $p$-dimensional random vector $Z_{i}\sim\mathcal{N}_{p}(\lambda 1_{p},I_{p})$, where $1_{p}$ is a vector of ones and $\lambda\neq 0$, and then taking $X_{i}:=Z_{i}/\|Z_{i}\|$. By symmetry, this strategy leads to the most central object always being equal to $\mu=\mathrm{sign}(\lambda)(1/\sqrt{p})1_{p}$ and the absolute value of the parameter $\lambda$ controls the spread of the points: the larger the value of $|\lambda|$, the more closely the points are concentrated around $\mu$. For the outliers we used $\lambda=-1$ and for the bulk $\lambda=5$, meaning that the two distributions have their most central points on the opposite sides of the hypersphere; see Figure C.1 in Appendix C for an illustration.
@@
The simulation thus has three parameters, $\varepsilon$, $p$ and $n$, which are set to the same values as in the earlier simulation. As shown in Figure 3 and its relative version presented in Appendix C, the performance of each metric depth function on the hypersphere dataset is very similar to its performance on the correlation data. We have omitted the timing results for the sphere simulation, as they were visually almost identical to those in Figure 2.
We also conducted an additional simulation in the same setting with $\varepsilon=0.10$, $n=100,200$, and $p=50,100,200,400,800,1600$ to investigate the effect of increasing the data dimensionality $p$ on the estimate. The average estimation errors over 100 replicates are plotted in Figure 4. The results indicate that: (i) MSD, MLD, MOD2, MOD3 yield almost identical results and, as expected, their performance improves as $n$ increases. (ii) MHD is worse than the other four and changing $n$ does not appear to affect it. The standard half-space depth is known to degenerate in higher dimensions [11] and it is therefore possible that MHD suffers from the same issue. A non-trivial solution could be to regularize its computation in a suitable way. (iii) For all methods, the estimation error increases with $p$. From the viewpoint of efficiency, for example $n=100$ and $p=500$ yields roughly the same accuracy as $n=200$ and $p=650$. (iv) As a baseline, we also considered an estimator which picks a random object from the sample and uses it as the estimate of the location. Regardless of $p$, this leads to estimation errors of $0.413$ and $0.4078$ for $n=100$ and $n=200$, respectively, showing that all estimators improve considerably upon this simple benchmark.
@@ 4.2 Out-of-sample optimization
The in-sample optimization studied in the previous section might come across as a naive approach, but in many metric spaces it is the best one can do, the lack of Euclidean structure preventing the use of standard optimization algorithms. However, despite being non-Euclidean, some metric spaces admit coordinate representations/encodings in Euclidean spaces. Such spaces include, e.g., unit spheres (stereographic projection), the positive definite manifold (Cholesky decomposition) and the space of $L_{2}$-functions (Karhunen-Loève expansion, assuming we truncate it). In these cases, Euclidean non-linear optimization can be used to find the deepest point.
To formalize the earlier, assume that there exists a bijective map $f:\mathcal{X}\to\mathcal{S}$ where $\mathcal{S}\subseteq\mathbb{R}^{q}$ for some $q$. Then, the problem of finding the deepest object of a sample $X_{1},\ldots,X_{n}\in\mathcal{X}$ can be formulated as
@@
where $D_{n}$ is some sample metric depth function and $v$ is an Euclidean vector. After finding the optimizer $v_{0}$, it can be mapped to $\mathcal{X}$ simply as $f(v_{0})$. Since the map $v\mapsto D_{n}(f^{-1}(v))$ is highly non-linear, we propose using bounded Nelder-Mead (NMKB) and Limited-memory BFGS (L-BFGS-B) for the optimization. NMKB is a robust and derivative-free optimizer, which makes it generally well-suited for non-smooth functions like metric depths. However, it can be slow in high-dimensional problems. On the other hand, L-BFGS-B uses approximated gradients to find the optimum solutions and stores only the most recent steps during the optimization process, making it faster and more memory-efficient. See [4] for more details.
@@
In this simulation, the matrix dimension $p$ was set to 5, and the contamination proportion $\varepsilon$ was fixed to 10%. The sample size, as in previous simulations in Sections 4.1.1 and 4.1.2, was taken to be $n=10,20,30,40,50,60$. A new parameter introduced in this simulation was the number of principal components, determined by a threshold $\texttt{TSH}\in(0,1]$, representing the minimum proportion of total variance to be explained by the chosen principal components. Thus, the number of principal components is $r=\max(2,\{k:\sum_{i=1}^{k}\text{ExplainedVariance}_{i}\geq\text{TSH}\})$. In this simulation TSH was set to be $0.9$. The whole simulation was repeated 500 times using both MHD and MOD3. As further competitors we took the MHD-based in-sample and out-of-sample estimators in the R-package MHD [8]; see also [7]. The latter is based on Riemannian optimization of the depth function and denoted by DLP hereafter. Finally, also the in-sample version of MOD3 was included.
Figure 5: Comparing the performance of different optimization methods on MOD3 and MHD. DLP refers to the out-of sample method. Left: estimation error of 7 different cases. L-FBSG-B yielded exactly the same result on MOD3 as NMKB (light red line). Right: Estimation Time. The scale of the y-axis is logarithmic.
As observed in the results in Figure 5 (left), the use of the two introduced optimization algorithms (NMKB and L-BFGS-B) combined with dimension reduction has significantly contributed to reducing the estimation error, compared to using the in-sample estimators indicated by the solid lines. We further observe that NMKB has the best performance, surpassing the other two optimization strategies, DLP and L-BFGS-B. Moreover, MOD3 exhibits lower estimation error than MHD in both out-of-sample and in-sample cases. However, as shown in the Figure 5 (right), MOD3 involves a higher computational cost compared to MHD. Additionally, despite the substantial computational burden imposed by using the NMKB algorithm, we also observe that the MOD3 in-sample estimator tends to follow a similar trend to its optimization counterparts.
@@ 5 Real data example
In this section, we evaluate the performances of the proposed depth functions using real data. The secondary aim of this experiment is to demonstrate how the depth functions, whose analytically complex form prevents theoretical inferential results, can still be used for statistical inference via relying on permutation and rank tests. Similar depth-based tests have also been considered earlier in the context of complex and high-dimensional data. For instance, [24] proposed several robust two-sample permutation tests for imaging data based on multivariate volume depth. Moreover, a depth-based global two-sample envelope test was developed in [23] for functional and image data, and [7] used depth-based rank testing proposed in [5] to compare samples of functional connectivity matrices.
For this experiment, the Age_pyramids_2014 data from R-package HistDAWass were chosen. This dataset contains histogram-valued data representing the age distribution of the population in each country worldwide, reported for both sexes combined as well as separately for males and females. We model the histograms as objects in the Wasserstein space $(\mathcal{X},d)$ where $d$ is the $L_{2}$-Wasserstein distance, implemented in the HistDAWass-package.
@@
The current analysis was conducted on a subset of data that includes both sexes across all African and European countries, accounting for a total of 95 countries. Based on the geographic locations of the two continents, we hypothesize that their age distributions differ. This is visually confirmed in Figure 6, where the deepest age distribution, w.r.t. MOD3, among African countries (left) shows a different structure from the corresponding country in Europe (right). Thus, our research question is: Is there a significant difference in the age distribution between these two groups? To study this, we next conduct a permutation test to assess whether the difference between the deepest histograms in the two continents is statistically significant. Letting $f_{1},f_{2}\in\mathcal{X}$ denote the deepest in-sample histograms of the two groups, we use the $L_{2}$-Wasserstein distance between them, $t=d(f_{1},f_{2})$, as the test statistic. To simulate the distribution of $t$ under the null hypothesis (i.e., assuming no difference between the two groups), we randomly permute the group labels of the $n=95$ histograms and compute the test statistic value. This reshuffling was done a total of 500 times, yielding $t^{*}_{1},\ldots,t^{*}_{500}$, using which the $p$-value of the test is computed as $\{t\leq t^{*}_{i}\}/500$. Because the experiment produces very small $p$-values, we rescaled the $y$-axis in Figure 7 to display $\log_{10}(p\text{-value})$ instead. Under this transformation, taller bars correspond to smaller $p$-values and therefore provide stronger evidence against the null hypothesis. All depth functions indicate a difference between the groups at the $0.05$ significance level; see the black bars in Figure 7 (left).
Figure 7: $p$-values of depth-based permutation and rank tests across all metric depth functions in the case of no contamination (left panel), 12 swaps (middle panel) and 16 swaps (right panel). Significance level of $\log_{10}(0.05)$ is indicated by blue dashed line. All $p$-values smaller than $10^{-16}$ were recorded as $10^{-16}$.
@@
We next take the previous result of a difference between the continents as a ground truth and continue the experiment by (a) contaminating the original data set, and (b) performing the same test for the contaminated data set and observing which of the depths still let us obtain the correct conclusion (group difference). This experiment therefore mimics the practical scenario where the data have been partially wrongly recorded. To perform the contamination, we randomly select $k$ countries and swap their labels (a European country is labeled an African country and vice versa), using two different choices, $k=12,16$. To reduce randomness in the choice of the swap, we repeat the simulation 10 times, using 500 iterations in the test each time and the resulting average $p$-values are shown as black bars in Figure 7 (middle and right). Under medium contamination (12 out of 95 countries mislabeled), all depths continue to yield the correct decision. Whereas, under high contamination (16 out of 95 countries mislabeled) both MHD and MOD2 identify the groups as indistinct. This is likely due to the fact that the algorithm used for MHD is approximate [7]. In addition, MOD2 is not a “true” depth function and therefore does not necessarily measure the centrality of its input object; see the discussion after Theorem 7. The most consistent performance across swaps is given by MOD3 and MSD, which likely relates to the fact that they both use numerical distance information in their computation, unlike MHD and MLD which only use indicators; see Section 3.2. The results also align with the simulation studies in Section 4.1 where MSD was almost always as good as MOD3.
Moreover, following [5, 7], we implemented depth-based rank tests using MLD, MHD, MSD, MOD2, and MOD3 for the same null hypothesis of no difference between the continents. The idea behind the tests is to inspect the ranks of the observation depths, which should show no group pattern under the null hypothesis. Two versions of the test described in [5] were considered: the K-statistic and the H-statistic. The K-statistic, essentially a Wilcoxon rank-sum test, uses pooled depth rankings but may lose power under location-shift alternatives (in the Euclidean case). The H-statistic is instead a more powerful alternative in which the depth rankings are computed separately relative to each group and then averaged across groups. Being depth-based, both statistics inherit all invariance properties of the underlying depth function. For both statistics, we used the asymptotic null distributions described in [5].
The $p$-values of the rank tests are shown in dark grey (K-statistic) and light grey (H-statistic) in Figure 7. As seen in the left panel, the depth-based rank test rejects the null hypothesis for the original data in all models except one: MHD under the K-type statistic does not lead to rejection at the conventional significance level. The corresponding numerical values are reported in Table C.1 in the appendix. This indicates that the permutation test provides more uniform sensitivity across depths, whereas the rank-based approach is more affected by the choices between the depths and the K- and H-type statistics. The results of the rank test under the swapping experiment in Figure 7 (middle and right) show that only MOD2 and MOD3, when combined with the H-type statistic, manage to reject the null under the strongest contamination (16 swaps). This is in line with the claim in [5], which reports that the K-type test is the less powerful of the two rank tests. Overall, among the depth functions considered, MOD3 appears to be the most stable, showing strong performance under both permutation and rank-based testing frameworks.